OpenStudy (anonymous):

5.00 g of hydrogen gas and 50.0g of oxygen gas are introduced into an otherwise empty 9.00L steel cylinder, and the hydrogen is ignitted by an electric spark. if the reaction product is gaseous water and the temperature of the cylinder is maintained at 35Celsius what is the final gas pressure inside the cylinder?

OpenStudy (jfraser):

You've got a limiting reactant case, where some reactant gas will be left over, and the product will also be a gas, so you need to find how many moles of water vapor is produced, AND how many moles of the excess reactant is left over. That total ## of moels gets plugged into PV=nRT to find total pressure, making sure to convert the temp to Kelvin

OpenStudy (anonymous):

ok so for the limiting reagent i got 2.48mol of H2O and the excess i got 3.13mol of H2O.....what is the total #of moles? do i add the 2 mol values i got?

OpenStudy (mos1635):

n=moles m=mass Mr=Relative molecular mass n=m/Mr so H2: n1=5/2=40/16=2.5 O2: n2=50/32=25/16=1.5625 2 H2 + O2 -> 2 H2O (oxygen is in execess) 2.5 1.5625 0 -2.5 -1.25 +2.5 0 0.3125 2.5 there for 0.3125+2.5=2.8125 mole of gas in cylinder P*V=n*R*T P=7.8925 atm